OpenStudy (anonymous):

Which of the following is the simplified form of x+5/x^2-25 times x-5/x^2-10x+25? a. 1/x^2-25 b. 1/x^2-10x+25 c. 1/x^2+10x+25 d. 1/x^2+25 Thanks!! :)

jimthompson5910 (jim_thompson5910):

(x+5)/(x^2-25) = (x+5)/((x+5)(x-5)) = 1/(x-5) So (x+5)/(x^2-25) = 1/(x-5) ------------------------------------------------------- (x-5)/(x^2-10x+25) = (x-5)/((x-5)(x-5)) = 1/(x-5) So (x-5)/(x^2-10x+25) = 1/(x-5) This means that the original expression then turns into (1/(x-5))*(1/(x-5)) which then multiplies out to 1/((x-5)(x-5)) 1/(x^2-10x+25) So the answer is choice B

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome